OpenStudy (anonymous):

Simplify i23? help please!

OpenStudy (anonymous):

i^23

OpenStudy (anonymous):

 @ksaichandra it looks like this

hartnn (hartnn):

u know what is 'i' ?

OpenStudy (anonymous):

do you mean i(23)

OpenStudy (anonymous):

oh i get it now but what is the meaning of the I

hartnn (hartnn):

@Kristy143SIO do u know the value of i ?

OpenStudy (anonymous):

i Stands for the square root of negative one i believe @hartnn @FweaBoyLive

hartnn (hartnn):

that is correct. so what is i^2 ?

OpenStudy (anonymous):

so i=-^23?

hartnn (hartnn):

\(\huge i=\sqrt{-1}\)

OpenStudy (anonymous):

@hartnn could it be a multiplication or division

OpenStudy (anonymous):

if its multiplication its just plain simnple and if it is division it has to be a decimal

hartnn (hartnn):

@Kristy143SIO what is i^2 and i^3 u know ?

OpenStudy (anonymous):

@hartnn I have no clue?

hartnn (hartnn):

ok, since \(\huge i=\sqrt {-1}, i^2 =-1, i^3=i.i^2=-i\\ \huge i^4=1\) did u get all these ???

OpenStudy (anonymous):

huh? @hartnn

hartnn (hartnn):

u knew i=sqrt{-1} isn't it ? so what do u get after squaring ? i^2 = ?

OpenStudy (anonymous):

-2? @hartnn

hartnn (hartnn):

no its not -2 \(i^2=\sqrt{-1}^2=-1\) ok ??

OpenStudy (anonymous):

ok thanks! gotcha! @hartnn

hartnn (hartnn):

so now could u find, i^3, i^4,..... ?

OpenStudy (anonymous):

they would both be -1? @hartnn

hartnn (hartnn):

nopes, see my previous comment, i have given their values..